3. OABC is a tetrahedron and OA = a, OB = b and OC = c. The points P and Q

are such that OA = AP and 2OB = BQ. The point M is the midpoint of P Q. Find

(i) AB, (ii) PQ, (iii) CQ, (iv) QM, (v) MB and (vi) OM in terms of a, b and c.


4. ABC is a triangle and P, Q are the midpoints of AB, AC respectively. If AB = 2x

and AC = 2y, express the vectors (i) BC, (ii) PQ, (iii) PC, (iv) BQ in terms of x

and y. What can you deduce about the directed line-segments BC and PQ?


5. ABC is a triangle. If D is the midpoint of AC, show that BA + BC = 2BD.

6. ABCD is a quadrilateral with AB equal and parallel to DC. Prove that AD is equal

and parallel to BC.


7. ABCD is a square and P, Q are the midpoints of BC, CD respectively. If AP = a

and AQ = b, find in terms of a and b, the directed line segments (i) AB, (ii) AD,

(iii) BD, (iv) AC.


8. ABC is a triangle and P is any point in BC. If PQ is the resultant of AP, PB, PC,

show that ABQC is a parallelogram, and Q is therefore a fixed point.


1
Expert's answer
2021-02-18T13:53:52-0500

Solution


Find

3.) (i) AB, (ii) PQ, (iii) CQ, (iv) QM, (v) MB and (vi) OM in terms of a, b and c.

Let "\vec{OA}=a,\ \vec{OB}=b,\ \vec{OC}=c"

Then

"\vec{OP}=2a,\ \vec{OQ}=3b\\
\vec{OA}+\vec{AB}=\vec{OB} \implies \vec{AB} =b-a\\
\vec{OP}+\vec{PQ}=\vec{OQ} \implies \vec{PQ} =3b-2a\\
\vec{OC}+\vec{CQ}=\vec{OQ} \implies \vec{CQ} =3b-c\\
\vec{QM}=-\frac{1}{2}\vec{PQ}=a-\frac{3}{2}b\\
\vec{OM}=\vec{OQ}+\vec{QM}=3b+a-\frac{3}{2}b=a+\frac{3}{2}b\\
\vec{OM}+\vec{MB}=\vec{OB} \implies \vec{MB}=b-(a+\frac{3}{2})=-a-\frac{1}{2}b"


4.) Express the vectors (i) BC, (ii) PQ, (iii) PC, (iv) BQ in terms of x

and y. What can you deduce about the directed line-segments BC and PQ?


So, as "2x + \vec{BC} = 2y", we get "\vec{BC} = 2y - 2x"

As "x + \vec{PQ} = y", we get "\vec{PQ} = y - x"

As "x + \vec{PC} = 2y", we get "\vec{PC} = 2y - x"

As "2x + \vec{BQ} = y", we get "\vec{BQ} = y - 2x"

Something we can deduce about the directed line-segments BC and PQ is that

"\vec{BC} = \vec{2PQ}" , so they are parallel.


5.) ABC is a triangle. If D is the midpoint of AC, show that BA + BC = 2BD.


Let "\vec{AB}=-x" and "\vec{BC}=y", Then "AC=AB+BC" which is equal to "-x+y". Then "BD=BA+AD =x+\frac{1}{2(-x+y)}=\frac{1}{2(x+y)}" which implies that "BA+BC=2BD= x+y= 2(\frac{1}{2x+y})" hence proved


6.) Prove that AD is equal and parallel to BC.


Join A and C (diagonal). In triangles ABC and ACD ∠BAC=∠ACD (alternate angles between parallels AB and DC); AC is common; AB=DC (given). So the triangles are congruent (SAS, two sides and included angle). Therefore ∠DAC=∠ACB, and DA=CB in length. But these are alternate angles (opposite sides of the diagonal), so DA is parallel to CB.


7.) Find in terms of a and b, the directed line segments (i) AB, (ii) AD, (iii) BD, (iv) AC.


Let "ABCD" be a square. Let equal sides of square = x.

Given that "P" and "Q" are middle points of "BC" and "CD" respectively


"BC=CD=x"

"BP=\frac{x}{2}\:and\:QD\:=\frac{x}{2}"


Given that "AP=a" and "AQ=b"

Now in right angled triangle "ABP"

By Pythagorous Theorem


"\left(AP\right)^2=\left(AB\right)^2+\left(BP\right)^2"

"a^2=x^2+\left(\frac{x}{2}\right)^2\\a^2=\frac{5x^2}{4}"


Now in right angled triangle ADQ

By Pythagorous Theorem


"\left(AQ\right)^2=\left(AD\right)^2+\left(DQ\right)^2\\
b^2=x^2+\frac{x^2}{4}"

"b^2=\frac{5x^2}{4}"


"a^2=b^2=\frac{5x^2}{4}\:or\:a=b=\frac{\sqrt{5x}}{2}"

"\:x=\frac{2a}{\sqrt{5}}\:or\:x=\frac{2b}{\sqrt{5}}"


"i.)\:AB=x=\frac{2a}{\sqrt{5}}\:or\:\frac{2b}{\sqrt{5}}""I.\:AB=x=\frac{2a}{\sqrt{5}}\:or\:\frac{2b}{\sqrt{5}}""I.\:AB=x=\frac{2a}{\sqrt{5}}\:or\:\frac{2b}{\sqrt{5}}"

"ii.)\:AD=x=\frac{2a}{\sqrt{5}}\:or\:\frac{2b}{\sqrt{5}}"

"iii.)\:BD"

"\left(BD\right)^2=\left(BC\right)^2+\left(CD\right)^2"

"=x^2+x^2"

"=2x^2"

"=2\times\:\frac{4a^2}{5}"

"BD=\sqrt{2}\times\frac{2a}{\sqrt{5}}\:or\:\sqrt{2}\times\frac{2b}{\sqrt{5}}"

"iv.)\:AC"

"\left(AC\right)^2=x^2+x^2=2x^2"

"\:AC=\frac{2\sqrt{2a}}{\sqrt{5}}\:or\:\frac{2\sqrt{2b}}{\sqrt{5}}"


8.) If "PQ" is the resultant of "AP, PB, PC", show that "ABQC" is a parallelogram, and "Q" is therefore a fixed point.











1) By adding all of the vectors we get the point "C'" and the vector "AC," which, when translated, results in the vector "PQ" at the last drawing.


2) "AP = C'Q" (the rule of translation), "BC' = PC" (when added vectors "BP" was translated into "CC'" vector). "C'Q" is parallel to "AP" so that the "C'" angle equals to the "P" angle. Thus, the C`BQ triangle equals the "PCA" triangle. Consequently, "BQ = AC," similarly, "AC = BQ, \ ABQC" is a parallelogram.